6a. Dissociation of the double stranded DNA helix into single coils is referred to as DNA melting. In this process, two strands of a DNA molecule can be separated by heat without breaking any covalent bonds. Every unique DNA molecule “melts” at a different temperature. Determine the order of the DNA sequences listed below according to relative melting temperatures (lowest Tm to highest Tm). Assume that they all begin as stable double-stranded DNA molecules. [3 pts] A. CTGAATCA B. CAGTGTCT C. TTAATCAT D. CCTGCGG 6b. Explain your answer for 6a.
Answer #1

DNA melting :-
• In this process no covalent bond is broken whereas hydrogen bonds are broken by temperature.
• DNA contains four bases- Guanine, Cytosine, Adenine, and Thymine. Hydrogen bonds found between these bases.
• Double hydrogen bonds present between adenine and thymine whereas triple hydrogen bonds present between cytosine and guanine.

• If there is more number of hydrogen bonds present in DNA, more temperature required for melting DNA.

Now, which DNA has more hydrogen bonds.
• A DNA with more number of cytosine and guanine has more hydrogen bonds (*because triple hydrogen bonds present between them which is more then double bonds between adenine and thymine).
So, A DNA strand with more Cytosine and guanine has higher melting temperature with compare to others.

Now in question :-
Number of cytosine and guanine in DNA-
(A) G+C = 3
(B) G+C= 4
(C) G+C = 1
(D) G+C = 6.
So, the sequence of melting point (from low to high) will be :-
C< A <B< D.